# Perception of Visual Content: Differences Between Humans and Foundation Models

### v1

@@ Introduction
However, since t-SNE is only a two-dimensional representation of the data, it would be useful to also look at the similarity scores, generated by the BLIP model’s cosine similarity function, between each pair of annotations across the dataset. This is shown in Table 1. These values give a more accurate and quantitative comparison of the annotation similarities. It can be seen that the ML object labels and the ML captions are the most similar pair of annotations (0.87), followed by ML captions and human labels (0.45), and ML object labels and human labels (0.42). It is important to note that when comparing human labels with ML-generated ones, the similarity scores are relatively low, regardless of the type of ML annotation.
The high similarity of both ML annotation types is a reasonable outcome. Although the annotations were generated by two different pre-trained ML models which were trained with different types of training data, i.e., labels versus image-text pairs, the training data both originated from the COCO dataset and are based on the same set of images. This means that there would be more overlap of objects between the two annotation categories. In contrast, the low similarity values between either ML annotations with the human labels is expected since there may be higher variety of words and phrases that were used by the human annotators. Moreover, when we compare the t-SNE visualisation and the similarity scores, it may seem inconsistent at first glance due to the ML objects and human labels pair having the lowest similarity score despite being seemingly furthest apart visually. However, since t-SNE is a dimensionality reduction technique, there may be features that were lost in the process, resulting in less accurate representations.
@@
To further explore the relationships of these similarity scores, a plot of pairwise relationships is presented in figure 3. The plot illustrates how similarity scores between two annotation categories correlate to another pair’s scores for a given image. The plot is colour-coded by region to allow further analysis of similarity scores by geographical location. The analysis reveals that there are no particular trends or patterns with respect to region. Moreover, when looking at the subplots for ML objects-ML captions pairs, i.e., subplots on first column/row, the similarity scores tend to be relatively high, on average, with the exception of several outliers. This aligns with the previous observation that ML objects and ML captions have the highest average similarity score (0.87). This pair also has the smallest variance out of all pairs, which is also evident when comparing the vertical spread of scores across the subplots in all rows, i.e., the scores of the ML objects-ML captions pair are not spread as much vertically when disregarding the outliers. In contrast, the ML objects-human labels pair, i.e., second row/column, appears to have the highest variance, which also aligns with the standard deviation of the mean similarity score observed in the previous experiment (0.1). It is also shown that the ML objects-ML captions pair has the same outliers as the ML objects-human labels pair (first row, second column), but not with the ML captions-human labels pair (first row, third column). Interestingly, the ML captions-human labels pair (third column/row) seems to have no outliers.
@@
The lack of discernible patterns in the way the scores are distributed across regions suggests that the images have little to no effect on how similar two sets of annotations will be. This could mean that the discriminative features in the images, or lack thereof, are detected equally by both humans and ML models. Therefore, none of the annotation sets describe images from a particular region any better or worse than the other. This also suggests that humans and machines may be similar in terms of how biased they are when perceiving images.
The low variance of similarity scores for the ML objects-ML captions pair suggests that for the majority of images, both annotations consistently contain similar objects, which - similar to the previous discussion on their high similarity scores - is expected due to the origin of the both models’ training data. Similarly, the high variance of scores for the ML objects-human labels pair aligns with expectations. This is due to, not only syntactic differences, i.e., objects vs longer phrases or sentences, but also differences in the way machines and humans interpret visual content, i.e., machines are more likely to capture easily recognisable objects and actions but may fail to capture more complex and nuanced observations, such as environment aesthetics or partially obstructed objects.
@@
Moreover, the fact that the ML objects-ML captions pair has the same outliers as the ML objects-human labels pair only (first row, second column) indicates that for those particular images, the ML object detection model may be detecting objects that do not appear in the corresponding ML captions or human labels, or failing to identify objects that do. This reasoning is consistent with the findings of Zhang et al. (2021), which suggest that deeper neural networks, such as ResNet - the architecture of our object detection model - may struggle more with difficult images. Perhaps there is obstruction or lack of objects in the image, leading to no or incorrect objects being detected in these outliers, resulting in the low similarity scores when compared to their respective ML Captions and human labels. The ML captioning model as well as human annotators have the advantage in that they are able to capture non-object elements, such as actions, e.g., “wash”, or environmental descriptions, e.g., “dirty”.
@@
RQ2: How do different combinations of annotations affect fairness in ML predictive models?
F1-Score (Mean ± Standard Deviation)
Annotations Used Africa Asia Europe The Americas Overall
ML Objects 0.43 ± 0.05 0.45 ± 0.06 0.30 ± 0.09 0.51 ± 0.10 0.42 ± 0.03
ML Captions 0.60 ± 0.05 0.32 ± 0.05 0.57 ± 0.02 0.38 ± 0.06 0.45 ± 0.02
Human Labels 0.62 ± 0.05 0.54 ± 0.13 0.57 ± 0.05 0.51 ± 0.11 0.56 ± 0.07
ML Objects & ML Captions 0.64 ± 0.07 0.50 ± 0.05 0.49 ± 0.06 0.46 ± 0.09 0.53 ± 0.04
ML Objects & Human Labels 0.57 ± 0.05 0.41 ± 0.05 0.57 ± 0.06 0.55 ± 0.10 0.51 ± 0.04
ML Captions & Human Labels 0.60 ± 0.09 0.38 ± 0.05 0.62 ± 0.07 0.51 ± 0.12 0.51 ± 0.03
ML Objects & ML Captions & Human Labels 0.65 ± 0.06 0.46 ± 0.04 0.51 ± 0.06 0.53 ± 0.11 0.53 ± 0.05
@@
Table 2: Comparison of the region classification model performance across different annotation types. The mean and standard deviation of the F1-Score were recorded over 10 independent runs, both for the overall model performance and for each class individually.
RMSE (Mean ± Standard Deviation)
Annotations Used Africa Asia Europe The Americas Overall
ML Obj 256.17 ± 88.2 1161.35 ± 116.23 2467.03 ± 250.75 1142.27 ± 83.42 1341.65 ± 85.56
ML Captions 179.69 ± 54.86 1456.47 ± 132.90 2164.40 ± 99.97 1815.89 ± 59.91 1564.56 ± 48.54
Human 228.35 ± 55.14 1657.22 ± 107.99 2675.59 ± 104.75 1433.51 ± 74.91 1633.25 ± 49.57
ML Obj & ML Captions 168.79 ± 39.57 1223.70 ± 137.91 2111.71 ± 138.83 1191.59 ± 136.96 1279.39 ± 57.66
ML Objects & Human 117.97 ± 11.04 1497.18 ± 86.39 2621.28 ± 219.83 1108.55 ± 85.18 1480.72 ± 78.63
ML Captions & Human 149.31 ± 16.17 1523.02 ± 108.62 2525.53 ± 186.66 1202.37 ± 72.59 1486.67 ± 86.93
Obj & Captions & Human 125.85 ± 8.27 1407.31 ± 54.97 2549.58 ± 135.91 1191.46 ± 82.55 1449.52 ± 45.33
@@
Table 3: Comparison of the income regression model performance across different annotation types. The mean and standard deviation of the RMSE were recorded over 10 independent runs, both for the overall model performance and for each class individually.
@@
Table 2 tabulates the F1-Scores of the region classification ensemble model across various combinations of the three annotation sets: ML Objects, ML Captions, and Human Labels. The mean and standard deviation of the F1-Scores across 10 independent runs were recorded, both for the overall model and for each class (region).
@@
It is discerned that the Human Labels resulted in the highest overall model performance (0.56). The models utilising combinations of the human and ML annotations exhibited poorer overall performance, followed by the ML Captions Model (0.45) and the ML Objects Model (0.42). Compared to other models, the Human Labels model has the smallest range of F1-Scores across all classes, ranging from 0.51 to 0.62. Moreover, when looking at per-class performance, it is revealed that each class best performs with different models. Africa (class) exhibits the highest performance with the model that uses all three annotation sets (0.65), Asia with the Human Labels Model (0.54), Europe with the ML Captions and Human Labels Model (0.62), and the Americas with the ML Objects and Human Labels Model (0.55). One notable observation is that regardless of overall or per-class performance, human labels are always used in the models - none of the ML-only models resulted in superior overall or per-class performance.
The superior overall performance of the Human Labels model is expected. As discussed by Zhang et al. (2021), humans are more likely to use contextual reasoning in their decision-making process. Although the paper talks about this in the context of humans segmenting images, this may similarly extend to annotation tasks, positively impacting the model’s performance in classifying regions. Perhaps it is the case that classifying regions in particular requires more than descriptions of objects and actions, both of which are what the ML models were trained on. This could likewise influence the fairer and more balanced class-level performance of the Human Labels model, as evidenced by the range of the F1-Scores across all classes. This may also account for why the best-performing models, both overall and class-level, are those that involve human labels.
While Zhang et al. (2021) claim that humans are not necessarily better than machines in choosing discriminative features for image segmentation, our classification model results highlight a different perspective. It is evident that the usage of human labels, whether on its own or in combination with another annotation set, leads to more accurate region classifications. From this, it can be deduced that humans are, in fact, better than machines at creating image annotations that contain these discriminative features in the context of region classification.
@@
Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the Human Labels model and the ML Objects model at $p$-value $<$ 0.001.
@@
Table 3 shows the Root Mean Square Error (RMSE) for the AdaBoost Regression algorithm used to predict income. Similar to the classification model, the experiment was conducted over 10 independent runs across different combinations of the three annotation sets. The mean and standard deviation of the RMSE values were obtained for the overall model and for each region.
It is apparent that, unlike the classification model, the regression model with ML Objects and ML Captions performed the best overall, indicated by its lowest RMSE value (1279.39). This model also has the smallest range of RMSE values across all classes, following the trend from the best-performing model in the previous experiment, ranging from 168.79 to 2111.71. On the class level, the best-performing models were the ML Objects, ML Object-Captions, and the ML Objects-Human models. However, like the classification model, it is revealed that the models responsible for the best overall and class-level performance are ones that contain ML Objects. Moreover, it is worth highlighting that Asia performs best with the ML Objects model, suggesting that the detected objects in the images from Asia are indicative of income.
The higher performance exhibited by the ML Objects and ML Captions Model is particularly unexpected, especially when contrasted with the trend found in the classification model. However, this observation aligns with the claim made by Zhang et al. (2021) on machines being better at selecting distinguishing features in the context of image segmentation. In our case of income regression, this also applies, although only for ML Objects rather than ML Captions. This could be attributed to certain objects in the images being associated with certain income levels, therefore improving the model’s ability to predict income. Further experiments would need to be conducted to explore this.
Furthermore, we compared the performance differences between the best and worst-performing models (overall), namely the ML Objects-ML Captions Model and the Human Model. The results of the Kolmogorov-Smirnov test showed that the difference is statistically significant ($p$ $<$ 0.001), suggesting that the ML Objects-ML Captions Model’s better performance is statistically reliable.
Figure 4 illustrates a scatter plot of the regression model trained using the ML Objects and ML Captions to predict income on the test dataset. The data points are colour-coded by continent. The blue dotted line represents the line of best fit and the red solid line represents the identity line.
The plot reveals that most of the samples whose incomes are underestimated are mostly countries from Asia, with some samples from the Americas and Europe. On the other hand, most of the overestimated samples are from the Americas, with several samples from Asian and Europe. It is also important to note that none of the African samples deviate significantly from the identity line.
The underestimation of samples from Asia may be attributed to the high number of samples within the lower to middle income range in the training set, resulting in the model skewing predictions for samples from Asia. Interestingly, when looking at the distribution of samples from the Americas, the income levels range from lower to middle income, as well, despite the overestimation by the model. This suggests that perhaps the features detected in the ML Objects and ML Captions for the Americas are not well-represented, causing inaccurate predictions. This is made worse due to the limited number of samples in the dataset, causing there to be less variety in income distribution for particular regions.
@@
Figure 4: Income Regression Model using ML Objects and ML Captions as Annotations Grouped by Continent
Limitations. First, the experiments were done on only one dataset due to the limited availability of image datasets that contain features relating to socio-economic background, such as geographical region and income. Performing the experiments on another similar image dataset would make the results more reliable. Second, the number of image samples in the dataset that have been annotated by humans is rather small. Due to time constraints and limited resources, doing further data collection to obtain more human annotations was not feasible given the scope of this investigation. Having more annotated images would result in better generalisability of the models. Third, due to the small size of the annotated data, under-sampling was not performed for the regression task - unlike the classification model which had under-sampling embedded in the algorithm - which meant that there may be data imbalance when considering number of samples in each region. This could potentially be the reason as to why the performance of some regions is notably better than others. Lastly, there is a risk of misclassification and fault intolerance in this work, as the small dataset could affect the generalisability of the model and, therefore, the trends observed.

### v2

@@ Introduction
Since t-SNE is only a two-dimensional representation of the data, it is useful to also look at the similarity scores, generated by the BLIP model’s cosine similarity function, between each pair of annotations across the dataset. This is shown in Table 1. These values give a more accurate and quantitative comparison of the annotation similarities. It can be seen that the ML Captions and Human Labels are the most similar pair of annotations (0.69), followed by ML Objects and ML Captions (0.50), and ML Objects and Human Labels (0.42). It is important to note that when comparing ML Objects to any of the other annotations, the similarity scores are relatively low. To further investigate length statistics per image, Human Labels averaged 7.18 words per worker, ML Captions averaged 8.62 words per caption, and ML Objects averaged 4.36 object labels per prediction. The length of ML Objects is determined by the number of objects in the predictions rather than the average number of words as ML Objects consists of a discrete list of items.
Figure 4: Pairwise relationships between similarity scores.
The higher similarity between ML Captions and Human Labels can be explained by several factors. First, both humans and ML captioning models can recognise not only objects but also actions in images, leading to overlapping features in their descriptions. Additionally, the likelihood of descriptive phrases appearing in both categories is much higher, which contributes to their high similarity. This is also further supported by the t-SNE visualisation which shows both of these annotations placed relatively close to each other in the same cluster. In contrast, ML Objects having a low similarity to both ML Captions and Human Labels is expected as ML Objects annotations are more likely to be syntactically different and may have a more limited range of vocabulary compared to the other annotation types. Moreover, the average length of the annotations shows that ML Captions and Human Labels have similar average lengths (8.62 and 7.18, respectively) compared to ML Objects (4.36). This further supports the interpretation that ML Captions are similar to Human Labels due to their sentence structures.
@@
To further explore the relationships of these similarity scores, a plot of pairwise relationships is presented in figure 4. The plot illustrates how similarity scores between two annotation categories correlate to another pair’s scores for a given image. The plot is colour-coded by region to allow further analysis of similarity scores by geographical location. The analysis reveals that most regions show no particular trends or patterns, with one exception: images from Europe tend to cluster in the higher similarity range (top right area) of the plots, indicating stronger agreement between all annotation types for European images. Moreover, when looking at the subplots for ML Captions-Human Labels pairs, i.e., subplots on the third column/row, the similarity scores tend to be relatively high, on average, with the exception of several outliers. This aligns with the previous observation that ML Captions and Human Labels have the highest average similarity score (0.69). It can also be seen that all three pairs show the same amount of spread of scores across all subplots, which correspond to the variance values of the mean similarity scores between each pair (0.12).
@@
The lack of discernible patterns in the way the scores are distributed across the Americas, Africa, and Asia suggests that images from these regions have little to no effect on how similar two sets of annotations will be. This could mean that the discriminative features in these regions’ images, or lack thereof, are detected equally by both humans and ML models. Therefore, none of the annotation sets describe images from these three regions any better or worse than the other. On the other hand, the cluster of European samples towards the top right of the plots suggests that European annotations from each group seem to share strong similarities with every other annotation type. This may mean that both humans and machines were able to perform more consistently across many European images compared to non-European ones. We hypothesise that this due to the fact that both humans and machine-based models are more familiar with Western societies images based on their past experience or training.
@@
Moreover, the fact that the ML Objects annotations have a relatively low similarity score indicates that the ML object detection model may be detecting objects that do not appear in the corresponding ML Captions or Human Labels, or failing to identify objects that do. This reasoning is consistent with the findings of Zhang et al. (2021), which suggest that deeper neural networks, such as ResNet - the architecture of our object detection model - may struggle more with difficult images. Perhaps there is obstruction or lack of objects in the image, leading to no or incorrect objects being detected in these outliers, resulting in the low similarity scores when compared to their respective ML Captions and Human Labels. The ML captioning model as well as human annotators have the advantage in that they are able to capture non-object elements, such as actions, e.g., “wash”, or environmental descriptions, e.g., “dirty”.
@@
RQ2: How do different combinations of annotations affect the performance and bias in ML predictive models?
Annotations Used F1-Score (Mean ± Standard Deviation)
Africa (471) Asia (471) Europe (472) The Americas (472) Overall (1,886)
ML Objects 0.46 ± 0.02 0.29 ± 0.02 0.41 ± 0.02 0.31 ± 0.03 0.37 ± 0.01
ML Captions 0.48 ± 0.02 0.35 ± 0.04 0.48 ± 0.02 0.34 ±0.04 0.41 ± 0.01
Human Labels 0.42 ± 0.02 0.32 ± 0.01 0.38 ± 0.02 0.29 ± 0.02 0.35 ± 0.01
ML Obj & ML Capt 0.42 ± 0.04 0.32 ± 0.03 0.47 ± 0.01 0.31 ± 0.03 0.38 ± 0.01
ML Obj & Human Labels 0.43 ± 0.01 0.32 ± 0.02 0.41 ± 0.02 0.29 ± 0.03 0.36 ± 0.01
ML Capt & Human Labels 0.42 ± 0.02 0.30 ± 0.02 0.40 ± 0.03 0.31 ± 0.02 0.36 ± 0.01
ML Obj, ML Capt, Human 0.50 ± 0.01 0.31 ± 0.02 0.44 ± 0.02 0.30 ± 0.02 0.39 ± 0.01
@@
Table 2: Comparison of the region classification model performance across different annotation types grouped by region (target variable), along with the number of samples per region and the overall total. Over 10 independent runs, the mean and standard deviation of the F1-Scores were recorded, both for the overall performance and each class (region) individually.
Annotations Used F1-Score (Mean ± Standard Deviation)
Drinking
Water
Drying
Clothes
Front
Doors
Hand
Washing
Kitchens
Living
Rooms
Places
for
Dinner
Washing
Clothes
Overall
(1,886)
MLO 0.35 ± 0.03 0.29 ± 0.07 0.28 ± 0.03 0.39 ± 0.03 0.44 ± 0.04 0.30 ± 0.03 0.40 ± 0.04 0.40 ± 0.04 0.37 ± 0.01
MLC 0.42 ± 0.04 0.38 ± 0.07 0.40 ± 0.05 0.50 ± 0.03 0.39 ± 0.04 0.37 ± 0.03 0.37 ± 0.04 0.40 ± 0.06 0.41 ± 0.01
H 0.32 ± 0.05 0.29 ± 0.04 0.18 ± 0.02 0.41 ± 0.02 0.45 ± 0.02 0.34 ± 0.03 0.28 ± 0.03 0.43 ± 0.02 0.35 ± 0.01
MLO & MLC 0.38 ± 0.04 0.45 ± 0.04 0.22 ± 0.03 0.48 ± 0.05 0.48 ± 0.03 0.27 ± 0.03 0.34 ± 0.02 0.39 ± 0.07 0.38 ± 0.01
MLO & H 0.38 ± 0.03 0.19 ± 0.03 0.30 ± 0.03 0.30 ± 0.02 0.40 ± 0.03 0.38 ± 0.02 0.43 ± 0.03 0.37 ± 0.03 0.36 ± 0.01
MLC & H 0.38 ± 0.02 0.45 ± 0.05 0.23 ± 0.03 0.38 ± 0.03 0.47 ± 0.03 0.20 ± 0.02 0.30 ± 0.04 0.43 ± 0.03 0.36 ± 0.01
MLO, MLC, H 0.32 ± 0.03 0.40 ± 0.04 0.29 ± 0.03 0.42 ± 0.02 0.45 ± 0.02 0.33 ± 0.02 0.39 ± 0.03 0.43 ± 0.04 0.39 ± 0.01
@@
Table 3: Comparison of the region classification model performance across different annotation types grouped by image category (image category was not included in training), along with the number of samples per category and the overall total. Over 10 independent runs, the mean and standard deviation of the F1-Scores were recorded, both for the overall model performance and each image category individually.
Annotations Used RMSE (Mean ± Standard Deviation)
Africa (471) Asia (471) Europe (472) The Americas (472) Overall (1,886)
ML Objects 1099.26 ± 106.03 1578.72 ± 91.81 2935.26 ± 85.79 1742.49 ± 55.11 1979.1 ± 45.47
ML Captions 601.73 ± 25.31 1634.27 ± 12.86 2973.81 ± 57.28 1620.76 ± 26.25 1925.1 ± 26.13
Human Labels 641.82 ± 30.95 1408.57 ± 35.75 2880.65 ± 24.67 1600.72 ± 29.54 1842.57 ± 14.99
ML Obj & ML Capt 622.80 ± 63.79 1468.67 ± 26.75 2777.27 ± 83.50 1624.42 ± 54.18 1817.49 ± 37.13
ML Obj & Human Labels 581.71 ± 28.81 1515.85 ± 164.94 2895.28 ± 31.16 1620.09 ± 35.57 1870.01 ± 41.36
ML Capt & Human Labels 569.12 ± 12.92 1420.02 ± 31.07 3297.25 ± 145.16 1578.29 ± 35.61 2006.36 ± 58.58
ML Obj, ML Capt, Human 561.42 ± 29.05 1439.70 ± 40.32 2890.99 ± 53.85 1585.46 ± 41.64 1842.89 ± 30.83
Table 4: Comparison of the income regression model performance across different annotation types grouped by region (target variable), along with the number of samples per region and the overall total. Over 10 independent runs, the mean and standard deviation of the RMSE were recorded, both for the overall model performance and each class (region) individually.
Annotations Used RMSE
Drinking
Water
Drying
Clothes
Front
Doors
Hand
Washing
Kitchens
Living
Rooms
Places
for
Dinner
Washing
Clothes
Overall
(1,886)
MLO 1738.35 1699.04 3026.46 1418.68 2502.32 1602.33 1421.24 2030.24 1979.10
±59.40 ±119.83 ±75.53 ±87.05 ±66.97 ±71.54 ±119.54 ±73.80 ±45.47
MLC 1765.65 1241.21 2864.27 1197.36 2389.16 1709.97 1593.86 2112.19 1925.10
±34.95 ±44.87 ±44.27 ±35.92 ±14.50 ±232.10 ±26.17 ±49.54 ±26.13
H 1571.56 1342.43 2744.72 1363.50 2365.00 1483.01 1242.21 2143.17 1842.57
±23.31 ±70.87 ±40.38 ±42.24 ±56.15 ±55.17 ±35.88 ±41.70 ±14.99
MLO & MLC 1531.87 1191.23 2758.96 1376.72 2354.67 1497.12 1341.66 2001.87 1817.49
±54.58 ±67.30 ±54.13 ±56.35 ±127.14 ±76.62 ±46.02 ±54.91 ±37.13
MLO & H 1545.01 1399.87 2729.15 1343.64 2385.23 1515.60 1412.33 2175.34 1870.01
±48.62 ±$4$26.66 ±30.48 ±43.41 ±23.03 ±293.80 ±41.00 ±41.36
MLC & H 1590.08 1246.10 2771.16 1258.17 2423.63 2818.41 1337.43 2057.95 2006.36
±49.66 ±29.81 ±20.11 ±19.13 ±55.94 ±407.44 ±32.27 ±26.72 ±58.58
MLO, MLC, H 1536.57 1250.52 2761.17 1294.69 2373.87 1538.69 1327.09 2143.24 1842.89
±48.04 ±36.12 ±20.38 ±42.16 ±94.76 ±109.56 ±70.26 ±55.33 ±30.83
Table 5: Comparison of the income regression model performance across different annotation types grouped by image category (image category was not included in training), along with the number of samples per category and the overall total. Over 10 independent runs, the mean and standard deviation of the RMSE were recorded, both for the overall model performance and each image category individually.
@@
Table 2 shows the F1-Scores of the region classification ensemble model across various combinations of the three annotation sets: ML Objects, ML Captions, and Human Labels. Table 5 shows the same F1-Scores, but is grouped into the different image categories in the data. The mean and standard deviation of the F1-Scores across 10 independent runs were recorded.
@@
We observe that ML Captions resulted in the highest overall model performance (0.41). The model utilising all three annotation sets exhibited slightly poorer overall performance (0.39), followed by the ML Objects and ML Captions Model (0.38), the ML Objects Model (0.37), the models utilising combinations of the Human Labels with either of the other two ML-based annotations (both at 0.36), and the Humans Label Model (0.35). Moreover, when looking at per-class performance, we can see that the ML Captions Model performs best for Asia, Europe, and the Americas, at 0.35, 0.48, and 0.34, respectively, whereas the model utilising a combination of all three sets performs best for Africa, showing possible discrimination of ML Captions towards scenes based in the African continent. One notable observation is that regardless of overall or per-class performance, ML Captions are always used in the models, whether it would be as a stand-alone or in combination with other annotations.
When looking through the per-category performance, it is evident that although the ML Captions Model continues to be the best-performing for the Drinking Water, Front Doors, and Hand Washing categories, all seven other categories are better suited with the other annotation models. The model containing both ML Objects and ML Captions was the most effective for the Drying Clothes and Kitchens categories. ML Objects and Human Labels excelled in handling the Living Rooms and Places for Dinner categories. ML Captions and Human Labels worked best for Drying Clothes and Washing Clothes. Moreover, Washing Clothes also achieved optimal performance with the Human Labels Model and the all-annotations model. It is also important to note that the best-performing models in the Washing Clothes category are ones that contain Human Labels. It is also the only category to have the Human Labels Model as its best-performing model. Similarly, Drying Clothes, which also has more than one optimal model, achieve best performance with annotation sets that contain ML Captions.
The superior overall performance of the ML Captions Model is expected. As discussed by Zhang et al. (2021), humans are not necessarily better than machines in choosing discriminative features. Although the paper talks about this in the context of humans segmenting images, this may similarly extend to annotation tasks, positively impacting the model’s performance in classifying regions. It is evident that the usage of ML Captions on their own leads to more accurate region classifications for Asia, Europe, and the Americas. This suggests that images from these regions may have discriminative features, such as objects and actions, that are captured by the ML Captions and reflected in the model’s predictions. Although the ML Captions Model worked quite well for African samples as well (0.48), combining all annotations seemed to perform better for this region. This suggests that there may be certain features that the ML Captions Model fails to capture on its own. It may be that there are objects that were detected by the ML Objects annotations that were not detected by the ML Captions, as captions tend to convey the overall meaning of the image and overlook objects in the background or unrelated to the main theme of the image. Additionally, the fact that ML Captions are always used in the best-performing models shows how the features captured by the captions tend to be very indicative of what region the image is from.
The variation in optimal models across the image categories is understandable. Images from each of the categories may contain varying objects and/or actions that are associated with certain geographical regions, e.g., for theHand Washing category, images of buckets may appear more frequently in samples from Africa, whereas images of sinks may appear more in samples from Europe. This explains why certain annotations are better than others at capturing relationships between the objects or actions and the regions. The best-performing models for five of the categories (Drying Clothes, Kitchens, Living Rooms, Places for Dinner, and Washing Clothes) are the ones with a combination of at least two annotation sets. This may be due to how the images from these categories may have discriminative features that cannot be captured with just one annotation set.
@@
Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the ML Captions Model and the Human Labels Model at $p$-value $<$ $1.15\times 10^{-7}$. This statistically significant value suggests that the ML Captions Model’s better performance is statistically reliable.
@@
Table 4 shows the Root Mean Square Error (RMSE) for the AdaBoost Regression algorithm used to predict income grouped by region, while Table 5 shows the scores grouped by image category. Similar to the classification model, the experiment was conducted over 10 independent runs across different combinations of the three annotation sets. The mean and standard deviation of the RMSE values were obtained for the overall model and for each region.
It is apparent that, unlike the classification model, the regression model with ML Objects and ML Captions performed the best overall, indicated by its lowest RMSE value (1817.49). The best-performing models differ across all regions. The model containing all annotations was the most effective for African images (561.42), similar to the classification model, while the Asian images achieved the best results with the Human Labels Model (1408). Europe saw optimal performance from the ML Objects and ML Captions Model (2777.27), and the Americas was best handled by the ML Captions and Human Labels Model (1578.29). Moreover, it is worth highlighting that Asia is the only region where a stand-alone annotation set achieved the best performance.
@@
Figure 5: Income Regression Model using ML Objects and Captions as Annotations Grouped by Continent. The x-axis represents the ground truth values, plotted in log scale, while the y-axis represents the predicted values, plotted in linear scale.
Looking at the performance of the models by category, it is evident that the ML Objects-ML Captions Model is the most consistent and achieves the best performance across half of the categories, namely Drinking Water, Drying Clothes, Kitchens, and Washing Clothes. The ML Objects-Human Labels Model was the most effective for Front Doors. The Handwashing category was best handled by the ML Captions Model, while the Human Labels Model achieved the best results for Living Rooms and Places for Dinner.
The observation that the ML Objects and ML Captions Model achieves highest performance aligns with the claim made by Zhang et al. (2021) on machines being better at selecting distinguishing features in the context of image segmentation. In our case of income regression, this also applies, although only for ML Objects rather than ML Captions. This could be attributed to certain objects in the images being associated with certain income levels, therefore improving the model’s ability to predict income. Further experiments would need to be conducted to explore this. Moreover, similar to the classification model, the all-annotations model worked best for African images, likely driven by the need for the annotations to encompass various features, including objects and actions. The images from Asia and the Americas were best handled by the Human Labels Model and the ML Captions-Human Labels Model, respectively. This shows how images from both regions have features indicative of income that are most likely not object-based, rather based on actions or nuanced descriptions of the environment. The fact that Asia required only Human Labels for its optimal performance suggests that the discriminative features for income still require full human intervention, and that although the ML annotations can still capture them, they do not perform as well, even when combined with the Human Labels. In contrast, the European images achieved best performance with the ML Objects-ML Captions Model. This suggests that the combined ML-based annotations are better at capturing the defining features and relationships to predict income for European images. This indicates a potential for bias in the ML annotations, potentially stemming from the training data of the ML models, which may be more representative of European content. However, another possibility is that the human annotators coincidentally managed to find better discriminating features for income prediction when annotating images from Asia.
The best regression models for the action-based image categories are either the ML Captions Model or the ML Objects-ML Captions Model, indicating how the ML models are better at finding the defining features to predict income. It may be due to how the images contain more easily identifiable actions and objects that are indicative of income. On the other hand, image categories that are not action-based, such as Living Rooms and Places for Dinner are best handled by the Human Labels Model. This suggests that humans are better at identifying key features for income prediction when said features are likely to be based on nuanced descriptions.
Furthermore, we compared the performance differences between the best and worst-performing models (overall), namely the ML Objects-ML Captions Model and the ML Captions-Human Model. The results of the Kolmogorov-Smirnov test showed that the difference is statistically significant ($p$ $<$ 0.001), suggesting that the ML Objects-ML Captions Model’s better performance is statistically reliable.
Figure 5 illustrates a scatter plot of the regression model trained using the ML Objects and ML Captions to predict income on the test dataset. Each data point represents an individual image sample from the dataset, which corresponds to a household. The points include associated attributes, such as region (continent), country, actual income, and predicted income. They are also colour-coded by continent. The blue dotted line represents the line of best fit and the red solid line represents the identity line. The x-axis represents the ground truth of the income, which has been plotted in log scale for a clearer visualisation, whereas the y-axis represents the predicted income, plotted in its original value.
The plot reveals that the samples whose incomes are underestimated come from all regions, with most of the Europe samples clustering towards the higher income area in the x-axis and the Africa samples clustering towards the lower income area in the x-axis. On the other hand, there are very few overestimated samples, most of which are only from Africa.
The underestimation of samples may be attributed to the high number of samples within the lower to middle income range in the dataset. Interestingly, when looking at the distribution, samples from Africa tend to be closer to the red identity line, indicating a more accurate prediction, especially compared to Europe. This is expected as this is also reflected in the RMSE scores of the regression models in Table 4. The low variance of the RMSE values for African samples may be due to the features and relationships in the African samples captured in the ML Objects and ML Captions being more different from those in other regions, enabling the model to be more consistent in its predictions. On the other hand, the key features found in the annotations of the European samples may be more complex, resulting in more variability in predictions, as shown in the vertical spread of the European samples in the scatter plot, as well as the greater variance in the RMSE table.
Limitations. Our study only considers certain ML models and certain types of images. We thus cannot claim that our observations would generalise beyond the data and model considered.